1.) The adsorption of CO on activated carbon was followed experimentally at 0°C. At the given e es, the following que pressures, the following quantities of adsorbed gas were measured a. Determine whether the measurements conform to the Langmuir isotherm b. Calculate the adsorption-desorption equilibrium constant KA c. Calculate the volume corresponding to a full monolayer coverage. Co V [cm3]10.3 19.3 27.3 34.1 4045.5PLease show all work. Thank you.
Answer #1
P V P/V
133 10.3 12.91262
267 19.3 13.8342
400 27.3 14.65201
533 34.1 15.6305
667 40 16.675
800 45.5 17.58242
P/V= (P/Vm) + (1/KVm)
slope 1/ Vm    = 0.007
Vm      = 142.8571
intercept 1/KVm    = 11.93
K         = 0.000587
a) Follows Langmuir isotherm
b) KA 0.000587
c) Vm 1704.286